OpenStudy (anonymous):

Find all numbers x for which |x - 3| = 8

myininaya (myininaya):

really?

OpenStudy (anonymous):

there are harder ones coming after this one

myininaya (myininaya):

lol

OpenStudy (anonymous):

I just want to do a sanity check before I tackle the other ones

myininaya (myininaya):

you know what you should do?

OpenStudy (anonymous):

abs(x-3) = x-3 if x-3 > 0 else -(x-3)

myininaya (myininaya):

you should ask them to solve both of these |x-3| <-8 and |x-3|=-8

OpenStudy (anonymous):

x-3 = 8 x= 11 and -(x-3) = 8 -x+3 = 8 x = -5

myininaya (myininaya):

i bet you a lot of them would get it wrong

OpenStudy (anonymous):

lol that is the very next question

myininaya (myininaya):

i don't know how many times i have told me on here |something| does not have negative output so |f(x)|=-8 has no solution |f(x)|>0 is never less than 0

OpenStudy (anonymous):

so x-3 = 8 or x-3 = -8 so -5 <= x <= 11 is that right?

OpenStudy (anonymous):

lol I'm wrong

myininaya (myininaya):

or you still talking about the equation you gave?

OpenStudy (anonymous):

yes

myininaya (myininaya):

that one guy has it right

OpenStudy (anonymous):

yeah :-D

myininaya (myininaya):

did you mean for it to be an inequality?

OpenStudy (anonymous):

no he's right

myininaya (myininaya):

ok

myininaya (myininaya):

omg i'm so bored

myininaya (myininaya):

i'm gonna go

OpenStudy (anonymous):

+-11

OpenStudy (anonymous):

oh sorry i mean x=+-5